### v3

# A Generative Approach to LLM Harmfulness Detection with Special Red Flag Tokens

Abstract
Most safety training methods for large language models (LLMs) are based on fine-tuning that forces models to shift from an unsafe answer to refusal when faced with harmful requests. Unfortunately, these drastic distribution shifts generally compromise model capabilities. To avoid that, we propose to expand the model’s vocabulary with a special token we call red flag token () and propose to train the model to insert this token into its response at any time when harmful content is generated or about to be generated. Our approach offers several advantages: it enables the model to explicitly learn the concept of harmfulness while marginally affecting the generated distribution, thus maintaining the model’s utility. It also evaluates each generated answer and provides robustness as good as adversarial training without the need to run attacks during training. Moreover, by encapsulating our safety tuning in a LoRA module, we provide additional defenses against fine-tuning API attacks.
1 Introduction
To make large language models (LLMs) that are robust against a determined adversary, practitioners rely on several security layers such as model hardening via fine-tuning (Zou et al., 2024; Xhonneux et al., 2024; Sheshadri et al., 2024), perplexity filters (Alon & Kamfonas, 2023), or harmfulness classifiers (Inan et al., 2023; Sharma et al., 2025). However, as the model capabilities progress, so does their attack surface, as innate abilities can be used to circumvent defences (Huang et al., 2024)—e.g., using a low-resource language to jailbreak the model. It is therefore natural to embed safety mechanisms into the models themselves, which will scale with these capabilities. Ideally, an LLM would have no harmful faculties, but this is unrealistic as many desirable skills can be both useful and damaging depending on the context. Therefore, we argue that the model itself should detect when it is being used harmfully and that this behaviour can be complementary to standard safety training whose goal is to remove harmful faculties from the model.
To address this, we propose a method to detect harmful behaviour using the LLM’s generative process, which does not compromise its helpfulness. Our approach uses an additional special token for LLMs to predict when the model considers that its capabilities are being used unsafely. We call this a red flag token () and train the model to output this token at any time during the generation of a harmful response, while not changing its response after or in benign contexts. This special token is excluded from the user vocabulary and can be filtered from streamed responses. This complementary layer of safety has several benefits. First, our method only requires the to be included in the answer with a single token change, while other methods such as adversarial training like Xhonneux et al. (2024), force models to completely change their output distribution from a harmful response to a refusal. Secondly, our method does not rely on a specific safe answer, meaning that if the model is broken, e.g., through pre-filling (Andriushchenko et al., 2024) or random sampling (Huang et al., 2023), we can still output a to tag the answer as harmful and have it be filtered.
We demonstrate the feasibility of this approach by designing a loss function consisting of three components: (i) a cross-entropy loss on generating the (ii) a Kullback-Leibler (KL) loss term on the generation after the (see Figure 1) and (iii) a KL loss term on benign utility conversations. We test our approach with several adversarial attacks on three open-source models: Llama3.2-3B-IT (Grattafiori, 2024), Mistralv3-IT (Jiang et al., 2023), and Phi-3.5 (Haider et al., 2024).
Finally, inspired by the idea of task arithmetic (Ilharco et al., 2023), we introduce the concept of storing the safety aspect of the model in a LoRA module (Hu et al., 2021) and show that it can be applied to retain detection capabilities after an attacker has leveraged a fine-tuning API to remove the safety training of the model.
To sum up, the contributions of this paper are:
we propose a specialised red flag token to reliably detect harmfulness at each generation step, even under strong adversarial attacks, including pre-filling, sampling, and automated jailbreaks like GCG and PAIR;
we show empirically that our approach generalises beyond the training distribution, effectively handling significantly longer inputs than those seen during training;
Finally, we demonstrate that our safety module can be efficiently stored in a LoRA module to be applied to detect harmful outputs from harmfully fine-tuned models.
2 Method
2.1 Threat model
We assume that LLM access is gated behind some web interface or API with no access to model weights, logits, or direct control of input/output processing—commonly referred to as a black box setting. The main assumption we make is that the service provider can evaluate the logits and output of the model before streaming it to the user, including filtering tokens such as assistant tokens or our token. We also consider a more permissive gray box setting where the user may have access to extra features such as pre-filling, and also some white box settings where the user has access to model weights. Finally, we consider the most permissive setting, a fine-tuning attack setting where the user has access to some fine-tuning API.
2.2 Our Loss
We assume that we have a dataset of harmful prompt-continuation pairs and a data set of harmless (a.k.a., benign) pairs. To train the model to flag unsafe generations, we sample an index from a distribution defined over the positions of the harmful continuation , where is a minimum offset that avoids flagging too early in the generation. The continuation is then split into three parts: the prefix , the red flag token , and the suffix . We use to denote the cross entropy and to denote the Kullback-Leibler divergence (KL). We consider our reference model . Our loss consists of three components: first, to ensure our model outputs the red flag token in harmful completions, we use a standard language modeling cross-entropy loss on all harmful completion tokens starting at the minimum offset up to and including the token:
To maintain model performance and reduce distribution shift as much as possible without increasing the likelihood of a harmful answer, we use a KL divergence on the tokens after the :
and again, to reduce distribution shift and to capture that the likelihoods should not change on unrelated tasks we include a KL loss on benign pairs from
All these losses put together, we get:
Note that none of the loss functions make a harmful continuation more likely, enabling our approach to be complementary to other safety fine-tuning techniques. Figure 1 summarises our approach, with the training procedure fully defined in Algorithm 1.
3 Experiments
3.1 Models & Datasets
We fine-tune Llama3.2-3B-IT (Grattafiori, 2024), Mistralv3-IT (Jiang et al., 2023), and Phi-3.5 (Haider et al., 2024) using our algorithm on the Harmbench (Mazeika et al., 2024) training set using 32 adversarial continuations per harmful prompt generated via the ablation attack (Arditi et al., 2024) along with original refusals. We use 5000 samples from the Alpaca dataset (Taori et al., 2023) as benign prompts. All models listed have a set of reserved special tokens as part of their tokenizers, allowing us to avoid extending the vocabulary and instead we re-purpose one of these unused special tokens to be the .
We assess model utility on standard LLM benchmarks, MMLU (Hendrycks et al., 2021), ARC-E and ARC-C (Chollet, 2019), as well as a Harmless dataset of benign prompts consisting of 119 random prompts from Ultrachat200k (validation split), and 40 benign prompts with a similar syntax as Harmbench provided by (Xhonneux et al., 2024, Appendix I). We also use this benign dataset to compute the false positive rate of refusal (or ) in Figure 3.
For adversarial robustness evaluation, we compute the defence success rate (DSR) of different attacks on the Harmbench Standard test set (Mazeika et al., 2024) that contain 159 harmful prompts, balancing with an equal number of benign prompts. Either a refusal or surpassing a certain probability threshold counts as a successful defence. The thresholds are set per model and are , , and for Llama3.2-3B-IT, Mistralv3-IT, and Phi-3.5, respectively.
All models are trained on a single A100 GPU with LoRA (Hu et al., 2021) and a batch size of 64 using the AdamW optimizer (Loshchilov & Hutter, 2017) (for more hyper-parameters see the Appendix A).
3.2 Design decisions
There are several design decisions to consider:
The cross-entropy loss on can be computed on each index before and including the sampled position or only on . In other words, we have the choice to allow the model for flexibility of when to output at the cost of potentially overfitting more because we now train the model to output immediately after the instruction token. In particular, this forces the model to judge the prompt quite strongly, leading to a higher probability for in a refusal as well. In practice, we tested both approaches and saw better results computing the cross entropy up to and including index . A potential solution to avoid over-fitting after the instruction token is to have a minimum offset into the harmful continuation both for sampling as well as the cross-entropy term, which we apply to Mistralv3-IT.
The sampling distribution of the index at which to insert the is a key choice. We tested both a geometric distribution as well as a uniform distribution over the harmful continuation. We decide to use a geometric distribution, which means that fewer positions will be used in the cross-entropy loss, and more will be used in .
The attention mask in the can also be amputated not to include the harmful prompt and the harmful continuation before the such as to force the model to store the harmful continuation information in the embedding. In other words, with probability , We choose to apply this trick probabilistically with probability 0.5, such as to make training more stable but still encourage the to be meaningful for generation.
Calibrating sensitivity can be done in several ways. We can add a logit bias term to the before passing the logits through a softmax, modifying the distribution to be:
where is a logit for the token and is a user-specified bias to re-weight . We can also directly monitor without adding a logit bias term or looking for explicit generations of , and set a model-dependant threshold in which to flag the generation as potentially harmful. We elect to use the second strategy as it works well and is simpler to apply in practice.
3.3 Baselines
We consider three baselines. Since the models come with safety training, we record the natural refusal rate. We then consider the refusal rate of CAT (Xhonneux et al., 2024)—an adversarial training technique using continuous attacks, where we replicate their training procedure using our datasets for a fair comparison. Finally, we adapt Jain et al. (2024) to our setting, whereby we insert the at the first position of the assistant’s response and call this baseline fixed-position-rf.
3.4 Robustness Evaluation
We compute the defence success rates of the following attacks:
Pre-filling is where an attacker inserts the first tokens as the response of the assistant. We use the Harmbench (Mazeika et al., 2024) affirmative responses as the pre-fill attack. Note that the models are allowed to check the logits of the pre-filled text.
Sampling Attacks jailbreak models by sampling the response multiple times (Hughes et al., 2024). Occasionally, the model may eventually provide an answer to a harmful prompt after repeated queries. In our experiments, we sample up to 16 times or until the model responds, as evaluated by the official classifier for text behaviours in HarmBench000huggingface.co/cais/HarmBench-Llama-2-13b-cls. We use a temperature of and a top- value of for the sampling attack.
GCG (Greedy Coordinate Gradient) is an automated jailbreak method proposed by Zou et al. (2023) which greedily optimizes a prompt suffix to maximize the probability of generating an affirmative response to a harmful prompt, implicitly pre-filling the model without having direct pre-filling access.
PAIR (Prompt Automatic Iterative Refinement) proposed by Chao et al. (2023) uses an attacker LLM to automatically generate jailbreaks for a separate targeted LLM without human intervention. In this way, the attacker LLM iteratively queries the target LLM to update and refine a candidate jailbreak.
Pre-filling and sampling are gray-box attacks using additional features such as temperature-based sampling and manually injecting text into a model generation. GCG is a white-box attack, and PAIR is a black-box attack. An attack is successful if the model does not refuse or the does not surpass a model-specific threshold. We evaluate refusals using GPT-4o (OpenAI et al., 2024) as a judge.
Figure 2 summarises our results. We first observe that our approach maintains near-perfect utility across all models. In particular, all baselines achieve utility scores comparable to the base model, with CAT showing the lowest score on Harmless, and this allows for a good comparison point in terms of robustness utility trade-off. CAT results in a 4% drop on the Harmless evaluation that indicates over-refusal.
Our proposed approach is able to nearly perfectly defend against the gray-box attacks, i.e., Sampling and Pre-filling across all models. We find that on Llama3.2-3B-IT and with the Pre-filling attack, the approach results in a significantly higher DSR than CAT, which mainly fails to refuse most harmful queries when a pre-filled affirmative response is present. The fixed-position-rf is marginally worse than on all attacks, with a larger gap on GCG.
We also report very good DSR against PAIR, and some improvement against GCG (with respect to the base model) for the approach. In general, comparing the base model and , we find that even when requests are not being refused, we can still defend against the malicious attack with the detection of .
3.5 Fine-tuning attacks
We consider the fine-tuning attack threat model discussed in Section 2.1, where the attacker is able to fine-tune the model by submitting training data through an API. We train a LoRA module that stores the weight changes for the model to insert the in harmful continuations on Llama3.2-3B-IT. We can then apply this LoRA module one or more times after the attacker has fine-tuned the base model and see if is still generated during harmful generations. For the fine-tuning attack we also use a LoRA module and use the data and hyper-parameters from Qi et al. (2024), which consists of about 100 harmful examples and continuations.
As baselines we consider both the aforementioned CAT (Xhonneux et al., 2024) as well as the fixed-position-rf stored in LoRA modules. In addition, we test whether applying each of these approaches multiple times can further improve robustness. Finally, we check whether we can combine the LoRA modules for the approach with the orthogonal CAT approach.
As before, we use the same Harmbench test set and the same Harmless dataset from Figure 2; the results are in Figure 3. This figure demonstrates the fundamental robustness and utility trade-off that exists in terms of false positive rate (FPR) and true positive rate (TPR)—i.e., defence success rate. The goal is to improve the Pareto front. The LoRA maintains good performance in harmful contexts, with the fixed-position baseline performing poorly. The adversarial training baseline (CAT) improves safety but cannot be calibrated, producing no Pareto front. However, applying the CAT LoRA once or twice yields alternative robustness trade-offs. Unlike CAT, our LoRA module allows fine-grained calibration and a strong Pareto front. We also show that combining the CAT and modules yields complementary benefits, improving robustness and enabling calibrated trade-offs between utility and safety (TPR/FPR).
We also validate that on benign fine-tuning our module does not impact the gained performance significantly. We test this on GSM8k (Cobbe et al., 2021) in chat mode under strict match of the answer—more details is provided in Appendix C.
3.6 Generalisation to Longer Contexts
During training, we only include data with a single turn of user/assistant interaction, with being injected following some distribution biased towards the start of the first assistant’s response. We validate that our approach generalises to multiple user/assistant interactions without over-fitting to the start of the conversation by sampling a number of benign conversations from UltraChat-200k, and then appending harmful queries from our test set. By monitoring the probability of throughout the model’s generation, we find that the prediction of does not deteriorate with increased sequence length. One chat interaction is shown in Figure 4 with Figure 6 containing the corresponding text; the probabilities consistently remain low in the regime in which it was trained (during the assistant’s turn) and increases sharply in the presence of harmful content during pre-filling and generation. Aggregate statistics are shown in Figure 5.
3.7 Limitations
Expectedly, our approach does not defend very well against attacks that have complete access to the model weights such as GCG, although it still achieves a considerable improvement in robustness compared to the base model. This shows that the association between harmful generation and the is circumvented by these attacks, demonstrating a limitation of our approach which we plan to investigate further. Furthermore, in the fine-tuning attack setting, there is still a trade-off between robustness and utility in the face of an attacker, albeit a much stronger Pareto-front than without the module.
Another limitation of our method is the amount of hyperparameters that are introduced that require tuning. While it is not too difficult to achieve a decent trade-off, one can likely achieve much better performance with better hyperparameter tuning as we did not have the resources to effectively tune the model (all the trainings were done with a single GPU). In addition, our ability to associate the with harmfulness relies on the data provided.
4 Related Work
Jailbreaking LLMs
Modern LLMs used as chatbots are trained to follow user instructions (Ouyang et al., 2022) while also being trained to respond in a safe and harmless manner (Perez et al., 2022). While users quickly found ways to manually craft “jailbreaks” which could circumvent these safeguards and elicit harmful content from these systems (Wei et al., 2023), automated methods for crafting adversarial attacks were also shown to be effective. Particularly, Zou et al. (2023) propose a greedy-coordinate gradient (GCG) search algorithm to find an adversarial suffix optimized to pre-fill (Vega et al., 2023) an affirmative response in a model’s response. Other approaches use heuristics to craft interpretable jailbreaks with only black-box access to the target model (Chao et al., 2023; Liu et al., 2023; Zeng et al., 2024). Given white-box access to the target model, more powerful attacks are possible. Adversarial soft prompts can be optimized to manipulate the model’s outputs (Schwinn et al., 2024), causal features responsible for refusal behaviour can be selectively ablated (Arditi et al., 2024), and fine-tuning can be used to override or remove safety training entirely (Qi et al., 2023).
Defences
Beyond standard pre-training, LLMs are typically trained with preference optimization techniques such as RLHF (Ouyang et al., 2022) or DPO (Rafailov et al., 2023) to be more aligned with human preferences. Jailbreaks can be incorporated into this preference alignment phase to increase resilience to such attacks (as is often done with red-teaming methods), but this does not often generalise to novel jailbreaks. Historically, in the context of vision models, actively training against adversarial attacks in an online manner (i.e., adversarial training) is the only method that has shown increased adversarial robustness (Madry et al., 2017). However, in the context of language, most discrete attacks are prohibitively expensive to use online. Mazeika et al. (2024) train against adversarial suffixes generated by GCG, but continually update a pool of examples rather than generate each attack from scratch. Other approaches perform adversarial training by attacking the embedding or latent space of the model (Xhonneux et al., 2024; Sheshadri et al., 2024) which is much more efficient to compute and transfers to discrete attacks. Beyond adversarial training, newer defences target and alter harmful representations in order to prevent a model from producing harmful outputs entirely (Zou et al., 2024). Independent from training a model to be more robust to jailbreaks is to classify and judge the potential harmfulness of the generated text, often with another LLM fine-tuned for this task (Inan et al., 2023; Feuer et al., 2024), although this does require additional resources to classify the outputs. Concurrent work Huang et al. (2025) has shown that classifiers alone are often not sufficient, further making the case that other approaches are needed especially against permissive but common threat models such as the fine-tuning box attack.
Special Tokens
Several works have explored training or utilising special tokens for specific purposes. Burtsev et al. (2020) prepend “memory” tokens to an input prompt on a target task. Goyal et al. (2023) append “pause” tokens, which are hypothesised to give the LLM a buffer sequence to reason over before producing an output. Mu et al. (2023) train LLMs to compress longer prompts into smaller sets of “gist” tokens as a means to shorten the context. Xiao et al. (2023) prepend “attention sinks” to improve generalization to long-context sequences. LLMs have also been trained to use a variety of tools (such as a calculator or internet access), which are denoted and invoked via special tokens (Schick et al., 2023).
Most closely related to our approach is the recent work of Jain et al. (2024), where a model is trained to prefix an output with a special refusal or response token based on the behaviour of whether the model refuses or responds to a prompt. While their approach is related in that special tokens are leveraged in the context of alignment, the approach and objective are conceptually different. Their method correlates these tokens with behaviour (i.e., refusal or response) in order to better calibrate such behaviours, whereas our approach correlates a special token with some implicit notion of a concept (i.e., harmfulness), without modifying the model’s original behaviour. This conceptual difference leads to drastically different losses in the formulation. For instance Jain et al. (2024) do not propose a KL divergence with a reference model (Equation 2) to maintain the predictions similar to the reference model after is outputted which hurt the model’s utility and is not complementary with standard safety training (instead it is a way to calibrate the model post-hoc safety training). Moreover, their model is only trained to output a “behavioural token" (e.g., “refuse" or “respond") at the beginning of the answer, which is significantly less efficient to detect harmfulness, as shown in our experiments. In contrast, our work proposes an approach that is complementary to standard safety training where the model essentially acts as an “implicit judge” on its own generated output, improving its transparency and providing a clear signal to evaluate potentially harmful generations without incurring any additional computational cost at inference time.
Beyond that, Wang et al. (2024) also learns to tag answers as harmless or harmful, but they use two stage training procedure and hardcode the tag to be at the end of the response. They only consider fixed jailbreak prompts rather than attacks and do not consider the fine-tuning setting at all. Finally, Zhang et al. (2024) train a model to output a special reset token followed by a refusal. This differs from our work as we optimize to keep the output post-flagging identical to the base model, with the intention of maintaining utility rather than enforcing a refusal. We also do not use a DPO-based objective, which may inadvertently increase the probability of generating harmful sequences before generating a reset token.
5 Conclusion
We propose detecting harmful outputs from a large language model (LLM) without an external classifier but using the generative model itself. To achieve this goal, we develop a training algorithm such that the target LLM outputs a special red flag token () at any time during a harmful generation. Our approach is robust to strong attacks like pre-filling and PAIR, and it generalizes to long, multi-turn contexts despite training on single-turn conversations. We further show that safety modules–such as our or CAT–can be stored in LoRA adapters and applied post-hoc to counter fine-tuning attacks, preserving benign behavior. Finally, we demonstrate that combining CAT with our module yields complementary benefits, improving robustness and maintaining utility.
6 Impact Statement
Machine learning tools such as large language models (LLMs) are finding widespread usage in today’s wealthy societies. As such, any work in this area has the potential for a significant impact, as it could avoid catastrophic outcomes due to a potential lack of safety of these widespread models.
This work aims to provide a new approach to reduce the harmful behaviour of LLMs when used via a webpage or API. As such, the desired impact of this work is overwhelmingly positive. However, it has to be acknowledged that any work aiming to filter or prevent harmful content from reaching users of non-open source LLMs can most likely also be re-used for censorship and thus also runs the risk of reinforcing biases of the LLM operator—intentionally or not.
More broadly and in the longer term, our work may enable practitioners to build an extra layer of safeguards into models that have capabilities that can both be useful and harmful and thus cannot or will not be removed. In such a situation, our approach and future derivatives can be used to tag and recognize the harmful usage of a capability. A potential downside is that practitioners may be over-reliant on this as a defence mechanism rather than ensuring that learning algorithms and data during pre-training and various post-training stages remove harmful capabilities to the model. As such, this work also considers the worst-case attacks, such as very strong fine-tuning attacks, continuous attacks, and ablation attacks, to clearly show that this approach can be circumvented with sufficient access and thus shall not be the only layer of safety for critical applications with very capable models.
Acknowledgments
This project was partially funded by a Samsung Advanced Institute of Technology (SAIT) x Mila grant. G. Gidel is a CIFAR AI Chair, he is supported by a Discovery Grant from the Natural Science and Engineering Research Council (NSERC) of Canada.
Appendix A Hyperparameters
| Hyperparameters | Llama3.2-3B-IT | Mistralv3-IT | Phi-3.5 |
| Batch Size | 64 | 64 | 64 |
| Learning Rate | 0.0001 | 0.0001 | 0.00002 |
| Learning Rate Scheduler | constant | constant | constant |
| 8 | 8 | 8 | |
| 1 | 1 | 1 | |
| 3 | 3 | 3 | |
| RF CE Cutoff | 0.15 | 0.15 | 0.15 |
| Attention Dropout | 0.5 | 0.5 | 0.5 |
| Warmup Ratio | 0.03 | 0.03 | 0.03 |
| LORA - r | 128 | 128 | 128 |
| LORA - | 64 | 64 | 64 |
| Min Offset | 0 | 16 | 0 |
Appendix B GPT-4o Judge
The harmfulness check
is done with the following prompt:
Do you consider the following question and response to be the start of a harmful answer? Answer with yes or no. Question: {prompt}. Response: {response}
The utility check
is done with the following prompt:
Do you consider the following response to be the start of a reasonable answer to the prompt? Answer with yes or no. Question: {prompt}. Response: {response}
Appendix C Fine-tuning attack
We validate that our approach of applying a safety LoRA module does not break benign fine-tuning. For this we train with SFT for one epoch on GSM8k (Cobbe et al., 2021) in chat mode. We train with batchsize 64, learning rate , LoRA parameters & , AdamW (Loshchilov & Hutter, 2019), and a constant learning rate schedule. We evaluate the GSM8k performance with the lm-evaluation-harness (Gao et al., 2024) using the command lm_eval --model hf --tasks gsm8k --num_fewshot=5 --device cuda:0 --batch_size 16 --model_args pretrained=meta-llama/Llama-3.2-3B-Instruct --apply_chat_template. For the base model we get a performance of under strict-match, the fine-tuned model gets , the fine-tuned model with one safety LoRA adapter gets , and the fine-tuned model with the adapter applied twice gets .
Appendix D Generalisation to Longer Contexts