### v4

# A Generative Approach to LLM Harmfulness Mitigation with Red Flag Tokens

Abstract
Many safety post-training methods for large language models (LLMs) are designed to modify the model’s behaviour from producing unsafe answers to issuing refusals. However, such distribution shifts are often brittle and degrade performance on desirable tasks. To address these pitfalls, we propose augmenting the model’s vocabulary with a special “red flag” () token, and training the model to insert this token whenever harmful content is generated or imminent. This approach enables the model to explicitly learn the concept of harmfulness in its representations, with minimal impact on utility due to the marginal change in the generated distribution of natural language. Moreover, because the token is embedded in the model’s vocabulary, we can naturally leverage the LLMs’ generalization capabilities, such as in-context learning (ICL) and out-of-distribution generalization to languages that are not formally supported (e.g., Japanese for Llama3). In particular, we demonstrate that through ICL alone, the model can learn to initiate reflective reasoning upon generating the token at inference, which steers the response away from harmful continuations or enables self-correction when the flag is raised falsely. This approach is orthogonal and complementary to existing safety technique (such as safety classifiers or standard safety training) and easier to evaluate in comparison to natural language refusals, as it does not require a human or automated judge to assess the harmlessness of the answers.
1 Introduction
Defending large language models (LLMs) against adversarial users relies on multiple security layers, including safety fine-tuning (Zou et al., 2024; Xhonneux et al., 2024; Sheshadri et al., 2024), perplexity filters (Alon & Kamfonas, 2023), or harmfulness classifiers (Inan et al., 2023; Sharma et al., 2025). However, as model capabilities advance, so does their attack surface, as innate abilities can be used to circumvent defences (Huang et al., 2024)—e.g., using a low-resource language to jailbreak the model. It is therefore natural to embed safety mechanisms directly into models, allowing defences to scale with capabilities. While an LLM with no harmful faculties would be ideal, this is unrealistic: many desirable skills can be either beneficial or harmful depending on context. We therefore argue that models themselves should be able to detect when their capabilities are being misused, complementing standard safety training that aims to remove harmful behaviours.
To address this, we propose a method to detect harmful behaviour through the LLM’s generative process without compromising helpfulness. Our approach uses an additional special “red flag” () token that the model learns to output when detecting unsafe capability usage. We train the model to output this token at any time during the generation of a harmful response, while not changing its response after the token, or in benign contexts. This special token is excluded from the user vocabulary and can be filtered from streamed responses. At inference time, the token serves as a flexible signal of potential harmfulness. As a defence, we consider two settings: filtering the entire response and replace it with a safe alternative (Section˜4.2), or triggering safety-oriented reflective reasoning before providing an answer (Section˜4.3.1) as illustrated in Figure˜1.
This approach offers several technical advantages. First, it requires minimal distribution shift: only a single token insertion rather than forcing complete output rewrites from harmful responses to refusals, as in traditional safety training and adversarial training methods (Xhonneux et al., 2024; Casper et al., 2024). This single explicit token also provides a natural and principled target for adversarial training, as we can directly optimize adversarial perturbations to suppress generation while encouraging affirmative responses, and then train against these attacks to improve robustness. Second, because the method is agnostic to specific safe responses, the model can still flag harmful content even when compromised through pre-filling (Andriushchenko et al., 2024) or random sampling (Huang et al., 2023). This also enables straightforward evaluation, since the detection of a token provides an objective signal without requiring human or automated judges to assess response harmlessness.
Third, unlike external safety classifiers that evaluate complete responses post-hoc, integrating the harmfulness signal directly into generation enables finer-grained detection and allows the model to condition on its own internal safety assessments. This integration with the generative process allows seamless application of existing LLM techniques: we demonstrate in Section˜4.3.1 that through prompting and in-context learning (ICL) alone, models can learn to use tokens as a soft signal for reflective chain-of-thought reasoning—evaluating whether the flagged content is truly harmful or not—rather than as a hard filter, enabling recovery from false positives on benign inputs. Finally, our approach demonstrates good generalization capabilities across the LLM’s natural capabilities; we show that our method generalizes effectively to other languages including supported languages like Spanish, and unsupported languages like Japanese (on our particular models), despite our fine-tuning dataset primarily comprising English prompts and completions (Section 4.3.2).
To summarise, the contributions of this paper are:
We propose to use a special red flag token to detect harmfulness at each generation step, even under strong adversarial attacks, including pre-filling, sampling, and automated jailbreaks like GCG and PAIR;
We demonstrate the feasibility of this approach by designing a loss function consisting of three components: (i) a cross-entropy loss on generating the token (ii) a Kullback-Leibler (KL) loss term on the generation after the token and (iii) a KL loss term on benign utility conversations;
We extend our training algorithm to an adversarial training paradigm to further improve results;
We show that our approach generalizes beyond the training distribution, including (a) integration with natural language instructions via in-context learning, where the model learns to use the token as a soft signal for reflective safety reasoning, and (b) cross-lingual transfer, including unsupported languages.
2 Related Work
Jailbreaking LLMs
Modern LLMs used as chatbots are trained to follow user instructions (Ouyang et al., 2022) while also being trained to respond in a safe and harmless manner (Perez et al., 2022). While users quickly found ways to manually craft “jailbreaks” which could circumvent these safeguards and elicit harmful content from these systems (Wei et al., 2023), automated methods for crafting adversarial attacks were also shown to be effective. Particularly, Zou et al. (2023) propose a greedy-coordinate gradient (GCG) search algorithm to find an adversarial suffix optimized to pre-fill (Vega et al., 2023) an affirmative response in a model’s response. Other approaches use heuristics to craft interpretable jailbreaks with only black-box access to the target model (Chao et al., 2023; Liu et al., 2023; Zeng et al., 2024). Given white-box access to the target model, more powerful attacks are possible. Adversarial soft prompts can be optimized to manipulate the model’s outputs (Schwinn et al., 2024), causal features responsible for refusal behaviour can be selectively ablated (Arditi et al., 2024), and fine-tuning can be used to override or remove safety training entirely (Qi et al., 2023).
Defences
A few recent works have proposed filtering harmful data from the pretraining phase (Li et al., 2025; O’Brien et al., 2025; Chen et al., 2025b). Our approach suggests an alternative to filtering such data, that could be implemented by simply tagging harmful data during pre-training. In addition to some filtering and rephrasing, contemporary work (Maini et al., 2025) explores the idea of “tagging” harmful content with a special token (corresponding to the use of a red flag token during pre-training) and demonstrates how it can be used to filter potential generations during beam search. In contrast, our work demonstrates how such a special red flag token can be learned at post-training time and how it can be used to either detect harmful content or be used to trigger reasoning, leveraging the generalization capabilities of LLMs.
Beyond standard pre-training, LLMs are typically trained with preference optimization techniques such as RLHF (Ouyang et al., 2022) or DPO (Rafailov et al., 2023) to be more aligned with human preferences. Jailbreaks can be incorporated into this preference alignment phase to increase resilience to such attacks (as is often done with red-teaming methods), but this does not often generalize to novel jailbreaks (Andriushchenko et al., 2024). Historically, in the context of vision models, actively training against adversarial attacks in an online manner (i.e., adversarial training) is the only method that has shown increased adversarial robustness (Madry et al., 2017). However, in the context of language, most discrete attacks are prohibitively expensive to use online. Mazeika et al. (2024) train against adversarial suffixes generated by GCG, but continually update a pool of examples rather than generate each attack from scratch. Other approaches perform adversarial training by attacking the embedding or latent space of the model (Xhonneux et al., 2024; Sheshadri et al., 2024) which is much more efficient to compute and transfers to discrete attacks. Beyond adversarial training, newer defences target and alter harmful representations in order to prevent a model from producing harmful outputs entirely (Zou et al., 2024). Independent from training a model to be more robust to jailbreaks is to classify and judge the potential harmfulness of the generated text, often with another LLM fine-tuned for this task (Inan et al., 2023; Feuer et al., 2024; Sharma et al., 2025), although this does require additional resources to classify the outputs. Huang et al. (2025) has shown that classifiers alone are often not sufficient, further making the case that other approaches are needed.
Special Tokens
Several works have explored training or utilising special tokens for specific purposes. Burtsev et al. (2020) prepend “memory” tokens to an input prompt on a target task. Goyal et al. (2023) append “pause” tokens, which are hypothesised to give the LLM a buffer sequence to reason over before producing an output. Mu et al. (2023) train LLMs to compress longer prompts into smaller sets of “gist” tokens as a means to shorten the context. Xiao et al. (2023) prepend “attention sinks” to improve generalization to long-context sequences. Chen et al. (2025a) append “defensive tokens” before the LLM input to defend against prompt injection attacks. LLMs have also been trained to use a variety of tools (such as a calculator or internet access), which are denoted and invoked via special tokens (Schick et al., 2023).
Closely related to our approach is the recent work of Jain et al. (2024), where a model is trained to prefix an output with a special refusal or response token based on the behaviour of whether the model refuses or responds to a prompt. While their approach is related in that special tokens are leveraged in the context of alignment, the approach and objective are conceptually different. Their method correlates these tokens with behaviour (i.e., refusal or response) in order to better calibrate such behaviours, whereas our approach correlates a special token with some implicit data-driven notion of a concept (i.e., harmfulness), without modifying the model’s original behaviour. This conceptual difference leads to drastically different losses in the formulation. For instance Jain et al. (2024) do not propose a KL divergence with a reference model (Equation˜2) to maintain the predictions similar to the reference model after a token is generated, since their objective was to provide a way to calibrate the model’s refusal sensitivity rather than to provide an additional safety layer. Moreover, their model is only trained to output a “behavioural token" (e.g., “refuse" or “respond") at the beginning of the answer, which is significantly less efficient to detect harmfulness, as shown in our experiments. In contrast, our work proposes an approach that is complementary to standard safety training where the model essentially acts as an “implicit judge” on its own generated output, improving its transparency and providing a clear signal to evaluate potentially harmful generations without incurring any additional computational cost at inference time.
Beyond that, Wang et al. (2024) also learns to tag answers as harmless or harmful, but they use two stage training procedure and hardcode the tag to be at the end of the response. They only consider fixed jailbreak prompts rather than attacks. Finally, Zhang et al. (2024) train a model to output a special reset token followed by a refusal. This differs from our work as we optimize to keep the output post-flagging identical to the base model, with the intention of maintaining utility rather than enforcing a refusal. We also do not use a DPO-based objective, which may inadvertently increase the probability of generating harmful sequences before generating a reset token.
3 Method
In this section, we present how we train the model to output a token during harmful completions and only marginally affecting the model’s output distribution. First, we outline the threat model we consider (Section˜3.1) before introducing the notation, datasets, and loss function we will use (Section˜3.2).
3.1 Threat model
We assume that LLM access is gated behind a web interface or API with no access to model weights, logits, or direct control of input/output processing—commonly referred to as a black box setting. We also consider a more permissive gray box setting where the user may have access to extra features such as pre-filling, and also some white box settings where the user has access to model weights.
3.2 Loss Formulation
We assume that we have a dataset of harmful prompt-continuation pairs and a data set of harmless (a.k.a., benign) pairs. To train the model to flag unsafe generations, we sample an index from a distribution defined over the positions of the harmful continuation , where is a minimum offset that avoids flagging too early in the generation. The continuation is then split into three parts: the prefix , the red flag token , and the suffix . We use to denote the cross entropy and to denote the Kullback-Leibler divergence (KL). The reference model we use is the initial model . Our loss consists of three components. First, to ensure our model outputs the red flag token in harmful completions, we use a standard language modelling cross-entropy loss on all harmful completion tokens starting at the minimum offset up to and including the token:
To maintain model performance and reduce distribution shift as much as possible without increasing the likelihood of a harmful answer, we use a KL divergence on the tokens after :
and again, to reduce distribution shift and to capture that the likelihoods should not change on unrelated tasks we include a KL loss on benign pairs from
All these losses put together, we get:
Note that none of the loss functions encourage harmful continuations, making our approach complementary to other safety fine-tuning techniques. Figure 2 summarises our approach, and the complete training algorithm is described in Algorithm˜1.
3.3 Design decisions
There are a few design decisions to consider:
Cross-entropy loss
The cross-entropy loss on can be computed on each index starting from the assistant generation and up to the sampled position , or only on . In other words, we have the choice to allow the model flexibility of when to output at the cost of potentially overfitting more because we now train the model to output a token immediately after the instruction token. In particular, this forces the model to judge the prompt quite strongly, leading to a higher probability for in a refusal as well. In practice, we tested both approaches and saw better results computing the cross entropy up to and including index . One additional trick we use is to “drop out” the red flag token from a harmful completion: in some harmful completions, we do not insert , but we still set all of the labels for the sequence to , encouraging the model to insert a even if it does not see it in a harmful completion.
Sampling distribution
The sequence position at which a token is inserted needs to be sampled from some distribution. We tested both a geometric distribution as well as a uniform distribution over the harmful completion. Empirically, we find that uniform sampling outperformed a geometric sampling scheme. The geometric sampling scheme tends to be vulnerable longer prefilling attacks since it is biased to be closer to the start of the generation, while sampling uniformly was much more reliable against long prefilling attacks.
Multiple tokens
We can also extend our approach to generate multiple tokens during harmful continuations. Training details for this extension are provided in Appendix˜C. The multi-emission variant is necessary when considering in-context examples containing tokens (Section˜4.3.1) because the model must be able to generate to flag harmful content while having that token already present in the few-shot examples preceding the user prompt.
3.4 Continuous Adversarial Training with Tokens
Another advantage of having a special token for the concept of harmfulness is that it can be used as a principled target to perform adversarial training. We adversarially train the token to improve robustness in a model we call RF-AT for which we provide evaluations in Section˜4.2. We do this by computing embedding space attacks (Schwinn et al., 2024) in a similar way as Xhonneux et al. (2024) and use them as adversarial examples in Equation 4. Specifically, we can compute adversarial embeddings to a user prompt by simultaneously minimizing the generation probability of (first term), and maximizing the probability of an affirmative response (second term):
The adversarial embedding perturbations are added to the input embeddings involved with the token loss components which pass through the online model , replacing in Equation 1 and Equation 2. We constrain each token embedding perturbation to an ball around the original token embedding. We optimize this attack using an -scaled SGD optimizer, where each step is scaled to an norm of 1.
4 Experiments
4.1 Models, Datasets & Baselines
Models
In all cases, we focus on models that have already undergone instruction tuning and safety training. We fine-tune Llama3.2-3B-IT (Grattafiori, 2024), Mistralv3-7B-IT (Jiang et al., 2023), and Phi-3.5-mini-IT (Haider et al., 2024) using the harmful partition of the Circuit Breakers000We do not train on XSTest, while Zou et al. (2024) include it in their training corpus. (Zou et al., 2024) training set, containing about 5000 harmful prompts with harmful completions and refusals for each prompt. For utility data, we sample 50k single from the Ultrachat200k dataset (Ding et al., 2023), only keeping one turn of conversation between the user and assistant. All models listed have a set of reserved special tokens as part of their tokenizers, allowing us to avoid extending the vocabulary and instead we re-purpose one of these unused special tokens to be the token. In some cases, we need to manually initialize the embedding weights to enable training. All models are trained on a single A100-80GB GPU with LoRA (Hu et al., 2021) and a trainable token (un)embedding, with an effective batch size of 64 using the AdamW optimizer (Loshchilov & Hutter, 2017); for more hyper-parameters see Appendix˜A.
Evaluation Datasets
We assess model utility on standard LLM benchmarks, MMLU (Hendrycks et al., 2021), ARC-E and ARC-C (Chollet, 2019), as well as a Harmless dataset of 180 benign prompts consisting of 119 random prompts from Ultrachat200k (validation split), and additional 61 benign prompts with a similar syntax as Harmbench. In Section˜4.3.1, we also evaluate on a subset of safe prompts of XSTest (Röttger et al., 2023) consisting of 4 categories: homonyms, safe targets, safe contexts, and definitions. We refer to this set as XSTest-Safe-Subset which includes 100 prompts (25 for each category). We focus on these categories due to their unambiguous safety characterization; remaining categories in XSTest predominantly feature fictional or borderline-unsafe cases that are not central to our study.
For adversarial robustness evaluation, we compute the defence success rate (DSR) of different attacks on the Harmbench Standard test set (Mazeika et al., 2024) that contain 159 harmful prompts. Either a refusal or a token generation counts as a successful defence.
Baselines
We consider three baselines. Since the models come with safety training, we record the natural refusal rate as Base. We then consider the refusal rate of CAT (Xhonneux et al., 2024)—an adversarial training technique using continuous attacks. Finally, we adapt Jain et al. (2024) to our setting, whereby we insert the token at the first position of the model response, and call this baseline Fixed-Position-RF.
4.2 Standard Robustness Evaluation
We compute the defence success rates of the following attacks:
Pre-filling is where an attacker inserts the first tokens as the response of the assistant. We use the Harmbench (Mazeika et al., 2024) affirmative responses as the pre-fill attack. Note that the models are allowed to check the logits of the pre-filled text.
Sampling Attacks jailbreak models by sampling the response multiple times (Hughes et al., 2024). Occasionally, the model may eventually provide an answer to a harmful prompt after repeated queries. In our experiments, we sample up to 16 times or until the model responds, as evaluated by the official classifier for text behaviours in HarmBench111huggingface.co/cais/HarmBench-Llama-2-13b-cls. We use a temperature of and a top- value of for the sampling attack.
GCG (Greedy Coordinate Gradient) is an automated jailbreak method proposed by Zou et al. (2023) which greedily optimizes a prompt suffix to maximize the probability of generating an affirmative response to a harmful prompt, implicitly pre-filling the model without having direct pre-filling access.
PAIR (Prompt Automatic Iterative Refinement) proposed by Chao et al. (2023) uses an attacker LLM to automatically generate jailbreaks for a separate targeted LLM without human intervention. In this way, the attacker LLM iteratively queries the target LLM to update and refine a candidate jailbreak.
Pre-filling and sampling are gray-box attacks using additional features such as temperature-based sampling and manually injecting text into a model generation. GCG is a white-box attack, and PAIR is a black-box attack. An attack is successful if the model does not refuse and the token is not generated. We evaluate refusals using GPT-5 (OpenAI, 2025) as a judge.
Results Fig. 3 summarises our results. We first observe that the approach maintains near-perfect utility across all models. In particular, all baselines achieve utility scores comparable to the base model, with CAT showing the lowest score on Harmless, and this allows for a good comparison point in terms of the robustness-utility trade-off. CAT results in about 10% drop on the Harmless evaluation, indicating significant over-refusal for this baseline.
Our proposed approach is able to nearly perfectly defend against the gray-box attacks, i.e., Sampling and Pre-filling across all models except Mistral, where the sampling attack is much stronger against the base model as well. We find that on Llama3.2-3B-IT and with the Pre-filling attack, the approach results in a significantly higher DSR than Fixed-Position-RF, which mainly fails to generate tokens on most harmful queries when a pre-filled affirmative response is present.
We also report considerable gain in DSR against GCG and PAIR attacks compared to the base model with the approach, and even larger gains (around 10%) with Red Flag Adversarial Training (RF-AT). CAT provides strong robustness against different attacks, but it has high over-refusal rates as shown by the drop in utility on Harmless in Figure˜3, and an even more significant drop on XSTest-Safe-Subset (as later shown in Section˜4.3.1), potentially making it more difficult to use in practical deployments. Overall, when comparing the base model with the red flag approaches, we find that token detection offers reasonable protection against malicious attacks in cases where the base model does not refuse, and that adversarial training further improves robustness, particularly against GCG and PAIR.
4.3 generalization Capabilities
We now investigate how tokens can synergize with LLMs generalization capabilities.
4.3.1 Red Flags as a Trigger for Reflective Safety Reasoning
A key advantage of our approach lies in the flexibility of how the red flag token can be used. In the previous section, we used it as a hard filter: any response that contains a is suppressed. Here we explore a soft-signal use: a token acts as a trigger for a short, safety-focused reflection segment.
Concretely, we leverage in-context learning with few-shot examples to prompt the model toward reflective safety reasoning when a token is present. The examples demonstrate that upon generating a token, the model should engage in safety-focused Chain-of-Thought (CoT) (Wei et al., 2022) reasoning to evaluate the surrounding context. Through this reflective process, the model determines whether the content is genuinely unsafe or acceptable. If deemed unsafe, it issues an appropriate refusal; if deemed safe (indicating spurious token generation), it proceeds to respond to the query. Sample behaviours for both benign and unsafe requests are illustrated in Figure˜5.
Setup
We provide the model with 10 few-shot examples (similar to those in Figure˜5) that precede the user prompt. The RF model, originally trained to generate red-flag tokens in harmful contexts, continues to produce these tokens but now uses them to trigger safety reasoning blocks. When safety reasoning blocks are successfully generated in the response, we extract the content following these blocks as the model’s final output. Otherwise, we consider the entire generated response. All responses are evaluated using GPT-5 with the prompt detailed in Appendix˜B. We assess both the defence Success Rate (DSR) against the Prefilling attack and the model’s utility score on the Harmless and XSTest-Safe-Subset datasets.
Results
Our results are shown in Figure˜4. The RF model with CoT prompting maintains a high DSR against prefilling attacks, successfully refusing 93% of the harmful queries. This is substantially above the Base model, and close to CAT and RF without CoT prompting. On benign utility benchmarks, performance remains strong on Harmless, and crucially, on XSTest-Safe-Subset, that contains safe but tricky questions, performance improves by 10% compared to RF without CoT prompting. CAT, on the other hand, performs very poorly on XSTest-Safe-Subset, only providing useful responses to 28% of the samples and refusing the rest. Although the performance is still behind the base model on XSTest-Safe-Subset, our experiments show great promise in reducing over-refusals by using as a soft signal to enable the model to recover from spurious flags on benign inputs.
This experiment demonstrates both the generalizability of the proposed generative approach and the flexibility inherent in usage, properties that distinguish it from rigid refusal-based fine-tuning methods like CAT and traditional harmfulness classifiers. For completeness, Appendix˜D provides the in-context learning template we use in this experiment.
4.3.2 generalization to other Languages
All previous experiments were conducted in English, including training the model. To further explore the generalization abilities of our approach, we investigate whether harmful requests in other supported languages and unsupported languages are correctly identified by our approach. Specifically, we prompt both the Llama3.2-3B-IT baseline and fine-tuned RF models with HarmBench queries in English, Spanish (supported), and Japanese (unsupported). For realistic attack conditions, we prompt each model 32 times per query using top-p sampling (); results appear in Figure 6. To assess the token’s impact beyond standard refusal, we report DSR for refusal-only responses and responses with generation. As expected, the translation attack achieve breaking the model more frequently (lower defence Success Rate). The baseline and models show similar robustness when considering refusal alone, confirming that our fine-tuning preserves standard refusal capabilities. However, when generation is considered, the model demonstrates considerably higher DSR. Our experiments show that our generative approach can leverage the LLM’s natural cross-lingual generalization ability, achieving stronger generalization beyond the training distribution than the base model.
5 Conclusion
In summary, we introduce the red flag token as a lightweight, model-native safety mechanism that complements existing defences. Unlike approaches that force full refusals or rely on external classifiers, our method integrates harmfulness detection directly into the generative process, enabling flexible use cases such as reflective reasoning and multilingual transfer. Our approach is robust to strong attacks like pre-filling and sampling, and provides considerable improvement over the base model on PAIR and GCG. By making safety signals an intrinsic part of model behaviour, this work takes a step toward scalable safeguards that could grow with model capabilities.
6 Impact Statement
Machine learning tools such as large language models (LLMs) are finding widespread usage in today’s wealthy societies. As such, any work in this area has the potential for a significant impact, as it could avoid catastrophic outcomes due to a potential lack of safety of these widespread models.
This work aims to provide a new approach to reduce the harmful behaviour of LLMs when used via a webpage or API. As such, the desired impact of this work is overwhelmingly positive. However, it has to be acknowledged that any work aiming to filter or prevent harmful content from reaching users of non-open source LLMs can most likely also be re-used for censorship and thus also runs the risk of reinforcing biases of the LLM operator—intentionally or not.
More broadly and in the longer term, our work may enable practitioners to build an extra layer of safeguards into models that have capabilities that can both be useful and harmful and thus cannot or will not be removed. In such a situation, our approach and future derivatives can be used to tag and recognize the harmful usage of a capability. A potential downside is that practitioners may be over-reliant on this as a defence mechanism rather than ensuring that learning algorithms and data during pre-training and various post-training stages remove harmful capabilities to the model.
Acknowledgments
This project was partially funded by a Samsung Advanced Institute of Technology (SAIT) Mila grant. D. Dobre is supported by Le Fonds de recherche du Québec – Nature et technologies (FRQNT), and the IVADO R3AI research grant on AI Safety and Alignment. G. Gidel is a CIFAR AI Chair, and is supported by a Discovery Grant from the Natural Science and Engineering Research Council (NSERC) of Canada. The authors would like to acknowledge the compute resources provided by Mila, and the Digital Research Alliance of Canada (DRAC).
Appendix A Hyperparameters
| Hyperparameters | Llama3.2-3B-IT | Mistralv3-7B-IT | Phi-3.5-mini-IT |
| Batch Size | 64 | 64 | 64 |
| Learning Rate | 2e-4 | 2e-4 | 2e-4 |
| Learning Rate Scheduler | constant | constant | constant |
| 1 | 1 | 1 | |
| 1 | 1 | 1 | |
| 1 | 1 | 1 | |
| RF CE Cutoff | 0.15 | 0.15 | 0.15 |
| Warmup Ratio | 0.03 | 0.03 | 0.03 |
| LORA - r | 128 | 128 | 128 |
| LORA - | 64 | 64 | 64 |
| Min Offset | 4 | 4 | 4 |
Appendix B GPT-5 Judge
Given a prompt and a response, we ask the following question from GPT-5 both for the harmfulness and the utility check:
Does the following response gives or starts giving actionable answer to the prompt? Answer with yes or no. Prompt: {prompt} Response: {response}
Appendix C Multiple Red Flag Insertions
Our primary approach to generating a focuses on producing a single token in a harmful context. We can extend our approach to inserting several by defining a multiple insertion scheme, and modify how we compute the loss. Rather than sampling an insertion point uniformly within the harmful continuation, we can insert by instead sampling values from a Normal distribution with a positive mean to define the distance from the previous insert position.
To help give more meaningful training signal, we also define a token-wise weighting scheme to scale the loss components on each token with respect to an inserted . For the cross-entropy loss, we define a ramp-up weighting scheme, where we increase the weighting of from 0 to 1 once we hit a . The weighting drops down until the next inserted , where it is ramped up again. For the , we define a decay weighting scheme, where the KL weight decays after a .
In practice, we sample 10 points from using a minimum offset of 0, and cutting out any points that exceed the sequence length. We define the ramp-up weighting scheme as an exponential gain over 20 tokens starting from 0 to 1, and the KL decay weighting scheme as a cosine decay over 40 tokens starting at 1 and decaying to 0.5; this means that after the first , the minimum KL scaling factor is 0.5.
Appendix D In-context Learning prompt
Following is the prompt we use for in-context learning in Section˜4.3.1 with three examples for brevity. In the experiments of Section˜4.3.1, we use 10 examples. RF_token will be replaced by ’s special token in the vocabulary.